Grade 12Physical Chemistry

Henry's law constant for in water is at 298 K. Calculate the quantity of in 500 mL of soda water when packed under 2.5 atm pressure at 298 K.

Murtaza
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
According to Henry's law,
pCO2 = KHx
Here x is the mole fraction of carbon dioxide in the carbon dioxide water mixture.
pCO2 = 2.5 atm = 2.5 * 1.01325 * 105 Pa = 2.533125 * 105  Pa
x = mole fraction = 2.533125 * 105/ 1.67 * 108 = 0.00152
x = moles of CO2/(moles of CO2 + moles of H2O) ≈moles of CO2/moles of H2O
as mole of CO2are negligible in comparison to the number of moles of water
500 ml of water = 500g of water = 500/18 = 27.78 moles of water.
So, moles of CO2/27.78 = 0.00152 or moles of CO2 = 0.042 moles.
Hence the amount of CO2 in 500 ml of soda water is 0.042 * 44 = 1.848 g